# Depth-guided NeRF Training via Earth Mover’s Distance

### v1

@@ Abstract
Neural Radiance Fields (NeRFs) are trained to minimize the rendering loss of predicted viewpoints. However, the photometric loss often does not provide enough information to disambiguate between different possible geometries yielding the same image. Previous work has thus incorporated depth supervision during NeRF training, leveraging dense predictions from pre-trained depth networks as pseudo-ground truth. While these depth priors are assumed to be perfect once filtered for noise, in practice, their accuracy is more challenging to capture. This work proposes a novel approach to uncertainty in depth priors for NeRF supervision. Instead of using custom-trained depth or uncertainty priors, we use off-the-shelf pretrained diffusion models to predict depth and capture uncertainty during the denoising process. Because we know that depth priors are prone to errors, we propose to supervise the ray termination distance distribution with Earth Mover’s Distance instead of enforcing the rendered depth to replicate the depth prior exactly through L2-loss. Our depth-guided NeRF outperforms all baselines on standard depth metrics by a large margin while maintaining performance on photometric measures.
@@ 1 Introduction
To improve the robustness of a neural radiance field in indoor settings, which are usually geometrically complex, previous work [24, 25, 6, 21] has attempted to incorporate depth supervision during training. The idea is that forcing NeRF to have a better understanding of the underlying scene geometry will enhance image rendering from novel view points. However, a challenge is that dense depth ground truths are rarely available or noisy.
[…]
[width=0.89]figures/pull4.eps
@@
Fortunately, a line of work has shown that depth estimation networks have become accurate at predicting the layout of indoor environments [20, 2, 16, 10, 19] and they can be used as pseudo ground truth for NeRF training. But when used ineffectively, depth supervision not only marginally improves photometric errors, but even fails to improve the NeRF’s depth renderings, suggesting a deficient understanding of scene geometry. One plausible reason for this phenomena is that, although depth prediction with neural networks has improved dramatically, there is still ambiguity owed to the ill-defined nature of estimating depth from one view, as many geometric layouts can lead to the same image. Partial occlusion of objects, shading, and reflections further challenge monocular depth prediction when applied out-of-domain, leading to misinterpreted geometries.
@@
We argue that although monocular depth predictions can be highly inaccurate, some are useful—so it is crucial to avoid a strict adherence to the depth prior in NeRF’s predictions. Depth priors should be a suggestion. Therefore, we propose to supervise the ray termination distance of a NeRF with the Earth Mover’s Distance ($EMD$) as illustrated in Figure 1. This invites the NeRF to sample ray termination distances close to the depth prior, without directly competing for the final weighting of ray termination distances with the photometric loss. Unlike previous works, we do not employ a uniform $L_{2}$-loss that will, inevitably, enforce incorrect depth as well, such as the missed window blinds in Figure 1. Instead, Earth Mover’s Distance allows us to maintain useful information about the distribution of the ray termination distances, while avoiding restrictive assumptions such as normally distributed errors, unimodality, continuity, or non-zero probabilities such as KL-divergence.
@@ 2 Related Work
Depth estimation from a single view is inherently ambiguous but many works have achieved impressive depth accuracy [12, 33, 2, 10, 22, 30]. [2] describes a method for depth estimation that generalizes well to multi-domain data by bridging the gap between relative and metric depth estimation. [30] uses a large-scale dataset and canonical camera transformations on input views to resolve the same metric ambiguity in zero-shot. [16, 28, 29] demonstrate how rich representations from self-supervised pretraining enables strong performance in dense visual perception. Most recently, [10, 22] leverage image-conditioned denoising diffusion in the depth estimation pipeline. In our work, we capitalize on the robust uncertainty estimates readily available in generative approaches to depth prediction. We use the monocular depth estimation network DiffDP [10] to provide depth pseudo ground truths for NeRF.
@@
Thus, recent works [9, 32, 15, 6, 21, 25, 8] have tried to enhance NeRF when only sparse views are available. Many of these rely on data-driven priors or regularization to guide NeRF optimization. [32] equips NeRF with per-view image features from a CNN encoder that is trained on large multi-view datasets, enabling scene reconstruction from as little as one view in a single forward pass. [9] supervises NeRF by enforcing consistency amongst CLIP [18] representations of arbitrary views. [15] samples unobserved views and regularizes the geometry and appearance of rendered patches from these views.
@@
Amongst these, a particularly interesting line of work explores supervising NeRF construction with depth [21, 27, 11, 6, 25, 8, 24]. Notably, few of these works explicitly identify the inherent uncertainty in depth predictions. [6] supervises NeRF ray sampling with COLMAP-derived [23] depth and models noise with a Gaussian centered at the depth predictions. [21] train a custom network in-domain to identify uncertain areas. However, these uncertainty maps overfit to the training case, where high depth errors usually occur at the edges of objects, and does not generalize to out-of-domain scenarios. [17, 8] focus on object-centric applications. [17] employs predicted depth by using the difference between sample location and estimated depth as conditioning for the NeRF and for depth-guided ray sampling. [8] supervises NeRF optimization by relaxing hard constraints in depth supervision to softer, more robust local depth ranking constraints. Although this approach acknowledges potential noise in pretrained depth network estimates, it does not directly incorporate the model’s uncertainty in these estimates to guide depth supervision for NeRF. SCADE [25] leverages 20 ambiguity-aware depth proposals generated by a costly out-of-domain prior network and a space-carving loss with mode-seeking behavior to supervise ray termination distance in NeRF. We found that while some depth hypotheses can be accurate, many of these proposals may also be poor predictors of scene geometry. This finding challenges the utility of having multimodal depth proposals for resolving ambiguity (see Figure 9), when NeRF naturally captures multimodal ray termination distributions. DäRF [24] is most closely related to our work and, like our work, proposes to use a standard pretrained depth estimation network to supervise NeRF training. However, DäRF does not consider pixel-wise uncertainty and supervises the NeRF-rendered depth directly instead of guiding the ray termination distance distribution, which allows for more fine-grained supervision of the NeRF’s learned scene representation.
@@ 3 Method
An overview of our method is shown in Figure 2. In this section, we describe each of the three components of our pipeline (depth prior construction, depth-guided ray termination sampling, and the photometric-geometric loss balancing) in turn.
@@ 3.2 Depth Priors and Uncertainty
An example uncertainty map is depicted in Figure 3, where DiffDP misinterprets the geometry of the cabinet, but highlights the same area as uncertain.
[…]
[width=0.7]figures/unc.eps
@@ 3.3 $EMD$-Guided Sampling
We use Earth Mover’s distance, as it naturally lends itself to compare probability densities or discrete histograms. It does not require satisfaction of assumptions such as that for two distributions ${P}$ and ${Q}$, ${Q}(y)=0\implies{P}(y)=0,\ \forall y$, which KL divergence would require, for example. Earth Mover’s distance also does not require that $y_{1},y_{2},...y_{N}$ are unimodally distributed, or that the error between NeRF-sampled ray termination distances and depth prior is normally distributed. Earth Mover’s distance is therefore equipped to address complex uncertainties in our depth priors, such as those we empirically observe and that previous methods have not sufficiently captured. As EMD is not differentiable, we use Sinkhorn Divergence [7] to approximate the $L_{\text{EMD}}$ during training.
@@ 3.4 Loss Weighting
where $\lambda$ is a balancing weight and $\gamma$ controls the impact of the uncertainties. But opposed to Focal Loss, we provide less weight when we are unsure about a pixel, instead of increasing the weight to force the model to learn hard examples.
@@ 4 Results
In this section we evaluate our proposed method in detail and compare it to existing works. Our evaluation is divided into two sections. First we show results highlighting the strengths of our method. The second portion will explain the importance of our design choices.
@@ 4.1 Experimental Setup
We evaluate our method on three ScanNet [5] scenes as previously chosen by DDPrior, DäRF, and SCADE [21, 25, 24]. Each scene consists of 18-20 training images and 8 test images. To measure the generalization capability of our framework, we also evaluate on an additional dataset which we refer to as ScanNet+ containing additional scenes from ScanNet. Unless otherwise stated, all experiments are performed and averaged over the three standard ScanNet scenes. For evaluation, we use the standard photometric measures PSNR, SSIM, and LPIPS [34] as used in the original NeRF paper [14] and in all baselines we compare to. As we are especially interested in the ability to render accurate depths, we also adapt the depth metrics used in DäRF, namely relative error (Rel) and RMSE [24].
We compare our work to several baselines that use depth-supervision for NeRF training in indoor environments. The work most related to ours is DäRF [24], as it also leverage a frozen out-of-the-box pre-trained monocular depth prediction network to supervise NeRF training. SCADE is related to our work as well, but uses a custom-trained depth prior, and does not evaluate the depth accuracy of their method. As SCADE provides pre-trained weights, we evaluate their method on all depth metrics. Additional baselines are a standard NeRF [14], DS-NeRF [6], and DDPrior [21], where we use the results reported by DDPrior. We report results on DDPrior using an out-of-domain prior as trained by SCADE, to replicate our setting where the depth network is trained on a different dataset.
@@ 4.3 Reconstruction Quality
[…]
[width=0.9]figures/bad_example.eps
Figure 4: Good RGB rendering quality does not imply good geometric understanding. In this test set example, SCADE [25] accurately renders the image (PSNR and SSIM above average), while misinterpreting the geometry of the depicted scene (five-fold average RMSE). The model does not capture the cabinet below the desk (purple pixels in the depth map indicate far objects, while yellow means close), but is still able to render the correct colors at the location of the cabinet.
In this section we evaluate our method’s ability to render novel views, and, importantly, its ability to understand the geometric layout of a scene. We thus report both photometric and depth-based metrics. But first, we demonstrate that photometric and depth errors do not always go hand in hand. Observe in Figure 4, that a model can yield great photometric results, while misunderstanding the geometry (e.g. such as errors due to textureless regions that are common in indoor environments). The remainder of this section will investigate the performance of our model on both metric categories.
@@
[…]
! RGB-based metrics Depth-based metrics D PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ SqRel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$ DiffDP [10] - - - 0.100 0.032 0.261 0.128 NeRF [14] 19.03 0.670 0.398 - - 1.163 - DS-NeRF [6] 20.85 0.713 0.344 - - 0.447 - DDPrior [21] 20.96 0.737 0.294 - - 0.236 - DDPrior* [21] 19.29 0.695 0.368 - - 0.474 - SCADE** [25] 21.54 0.732 0.292 0.086 0.030 0.252 0.118 DäRF [24] 21.58 0.765 0.325 0.151 0.071 0.356 0.168 Ours 21.69 0.737 0.373 0.070 0.024 0.221 0.105 * Trained by SCADE authors. ** Depth metrics evaluated by us based on SCADE’s weights.
[…]
[width=]figures/qualitative_results-5.eps
@@
We compare our method on RGB-based and depth-based metrics in Table 1. Our method reduces all depth metrics of all baselines by at least $11\%$. Our model reduces the error of the most related work, DäRF, even by $54\%,66\%,38\%$, and $38\%$, respectively, on the four depth metrics. Our method also outperforms DDPrior [21] whose depth network was trained in-domain. Additionaly, our NeRF setup improves the initial depth estimates of the depth prior DiffDP. Compared to SCADE and DäRF, we perform on par on photometric measures, as each method wins on one metric.
Note that our method not only outperforms SCADE, but also avoids the need for a custom-trained depth prior. Training this prior requires evaluating the entire training set 20 times after each epoch to find the best hypothesis for each training example. Our training of SCADE’s cIMLE-based prior takes more than one week on four 24GB NVIDIA Titan RTX GPUs, despite performing the described procedure only every 10th epoch. Details are discussed in Section 4.4.
[…]
[width=]figures/qualitative_results_depth-1.eps
@@
[…]
0.5! PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$ SCADE [25] 20.57 0.687 0.369 0.396 1.032 DäRF [24] 22.01 0.751 0.319 0.745 1.765 Ours 22.29 0.722 0.391 0.156 0.456
We qualitatively compare the results in Figure 5. In the illustrated examples, our method leads to less severe artifacts, and crisper and more accurate edges in both RGB renderings and depth maps. Additional qualitative examples of depth maps are shown in Figure 6. Our model more accurately reconstruct the depicted layouts, like the area underneath the piano in the second image, or the area under the table in the third image. Interestingly, DäRF’s depths are extremely crisp. However, the quantitative evaluation in Table 1 confirms that DäRF learns an inaccurate representation of the geometry. Its strong $L_{2}$-loss enforces the depth maps to look like their prior, although they do not correctly represent the scene.
To demonstrate the robustness of our method we report results on ScanNet+ that were not previously used by the baselines. We observe in Table 2, that our method is able to reduce the RMSE of the baselines by at least $56\%$, while producing comparable photometric results.
Additional qualitative results are shown in the supplementary material.

### v2

# https://anitarau.github.io/emd-nerf.github.io/

@@ Abstract
Neural Radiance Fields (NeRFs) are trained to minimize the rendering loss of predicted viewpoints. However, the photometric loss often does not provide enough information to disambiguate between different possible geometries yielding the same image. Previous work has thus incorporated depth supervision during NeRF training, leveraging dense predictions from pre-trained depth networks as pseudo-ground truth. While these depth priors are assumed to be perfect once filtered for noise, in practice, their accuracy is more challenging to capture. This work proposes a novel approach to uncertainty in depth priors for NeRF supervision. Instead of using custom-trained depth or uncertainty priors, we use off-the-shelf pretrained diffusion models to predict depth and capture uncertainty during the denoising process. Because we know that depth priors are prone to errors, we propose to supervise the ray termination distance distribution with Earth Mover’s Distance instead of enforcing the rendered depth to replicate the depth prior exactly through $L_{2}$-loss. Our depth-guided NeRF outperforms all baselines on standard depth metrics by a large margin while maintaining performance on photometric measures.
@@ 1 Introduction
To improve the robustness of NeRFs in complex indoor settings, previous work [25, 26, 7, 22] has incorporated depth supervision during training. The idea is that a better understanding of the underlying scene geometry will enhance image rendering. However, dense depth ground truths are rarely available.
@@
Fortunately, a line of work has shown that depth estimation networks have become accurate at predicting the layout of indoor environments [21, 2, 17, 11, 20] and they can be used as pseudo ground truth for NeRF training. But when used ineffectively, depth supervision can fail to improve the NeRF’s depth renderings, suggesting a deficient understanding of scene geometry. One reason is that, although depth prediction with neural networks has improved dramatically, there is still ambiguity owed to the ill-defined nature of estimating depth from one view, as many geometric layouts can lead to the same image. Partial occlusion of objects, shading, and reflections further challenge monocular depth prediction, leading to misinterpreted geometries.
@@
We argue that although monocular depth predictions can be highly inaccurate, some are useful—so it is crucial to avoid a strict adherence to the depth prior in NeRF’s predictions. Depth priors should be a suggestion. We propose to supervise the ray termination distance of a NeRF with the Earth Mover’s Distance ($EMD$) as illustrated in Figure 1. This invites the NeRF to sample ray termination distances close to the depth prior, without directly competing for the final weighting of ray termination distances with the photometric loss. Unlike previous works, we avoid $L_{2}$-losses that will, inevitably, enforce incorrect depth as well, such as the missed window blinds in Figure 1. Instead, EMD allows us to maintain useful information about the distribution of the ray termination distances, while avoiding restrictive assumptions such as normally distributed errors, unimodality, continuity, or non-zero probabilities such as KL-divergence.
@@ 2 Related Work
Depth estimation from a single view is inherently ambiguous but many works have achieved impressive depth accuracy [13, 34, 2, 11, 23, 31, 30]. [2] describes a method for depth estimation that generalizes well to multi-domain data by bridging the gap between relative and metric depth estimation. [31] uses a large-scale dataset and canonical camera transformations on input views to resolve the same metric ambiguity in zero-shot. [17, 29, 30] demonstrate how rich representations from self-supervised pretraining enables strong performance in dense visual perception. Most recently, [11, 23] leverage image-conditioned denoising diffusion in the depth estimation pipeline. In our work, we capitalize on the robust uncertainty estimates readily available in generative approaches to depth prediction. We use the monocular depth estimation network DiffDP  [^1] Both [11] and [22] are called DDP, so we refer to them as DiffDP and DDPrior. [11] to provide depth pseudo ground truths for NeRF.
@@
Thus, recent works [10, 33, 16, 7, 22, 26, 9] have enhanced NeRF when only sparse views are available. Many of these rely on data-driven priors or regularization to guide NeRF optimization. [33] equips NeRF with per-view image features from a CNN encoder that is trained on large multi-view datasets, enabling scene reconstruction from as little as one view in a single forward pass. [10] supervises NeRF by enforcing consistency amongst CLIP [19] representations of arbitrary views. [16] samples unobserved views and regularizes the geometry and appearance of rendered patches from these views.
@@
Amongst these, an interesting line of work explores supervising NeRF construction with depth [22, 28, 12, 7, 26, 9, 25]. Notably, few of these works explicitly identify the inherent uncertainty in depth predictions. [7] supervises NeRF ray sampling with COLMAP-derived [24] depth and models noise with a Gaussian centered at the depth predictions. [22] train a custom network in-domain to predict uncertain areas. However, the model overfits to the training case, where high depth errors occur at the edges of objects. [18, 9] focus on object-centric applications with large overlap between train images. [18] employs predicted depth by using the difference between sample location and estimated depth as conditioning for the NeRF and for depth-guided ray sampling. [9] supervises NeRF optimization by relaxing hard constraints in depth supervision to softer, more robust local depth ranking constraints. Although this approach acknowledges potential noise in pretrained depth network estimates, it does not directly incorporate uncertainty to guide depth supervision for NeRF. SCADE [26] leverages 20 ambiguity-aware depth proposals generated by a costly out-of-domain prior network and a space-carving loss with mode-seeking behavior to supervise ray termination distance in NeRF. We found that while some depth hypotheses can be accurate, many of these proposals may also be poor predictors of scene geometry. This finding challenges the utility of having multimodal depth proposals for resolving ambiguity (see Figure 9), when NeRF naturally captures multimodal ray termination distributions. DäRF [25] is most closely related to our work and, like our work, proposes to use a standard pretrained depth estimation network to supervise NeRF training. Yet, DäRF does not consider pixel-wise uncertainty and supervises the NeRF-rendered depth directly instead of guiding the ray termination distance distribution.
@@ 3 Method
An overview of our method is shown in Figure 2. In this section, we describe each of the three components of our pipeline (depth prior, EMD-guided ray termination sampling, and photometric-geometric loss balancing) in turn.
@@ 3.2 Depth Priors and Uncertainty
An example uncertainty map is depicted in Figure 3, where DiffDP misinterprets the geometry of the electrical box, but highlights the same area as uncertain.
@@ 3.3 $EMD$-Guided Sampling
We use Earth Mover’s distance, as it naturally lends itself to compare probability densities or discrete histograms. It does not require satisfaction of assumptions such as that for two distributions ${P}$ and ${Q}$, ${Q}(y)=0\implies{P}(y)=0,\ \forall y$, which KL divergence would require. Earth Mover’s distance also does not require that $y_{1},y_{2},...y_{N}$ are unimodally distributed, or that the error between NeRF-sampled ray termination distances and depth prior is normally distributed. Earth Mover’s distance is therefore equipped to address complex uncertainties in our depth priors, such as those we empirically observe and that previous methods have not sufficiently captured. As EMD is not differentiable, we use Sinkhorn Divergence [8] to approximate the $L_{\text{EMD}}$ during training.
@@ 3.4 Loss Weighting
where $\lambda$ is a balancing weight and $\gamma$ controls the impact of the uncertainties $u$. Opposed to Focal Loss which increases the weight of uncertain examples to force the model to learn hard cases, we apply less weight to uncertain pixels.
@@ 4 Results
In this section we evaluate the proposed method in detail, compare it to existing work, and explore the importance of our design choices.
@@ 4.1 Experimental Setup
We evaluate our method on three ScanNet [5] scenes as chosen by DDPrior, DäRF, and SCADE [22, 26, 25]. Each scene consists of 18-20 training images and 8 test images. To measure the generalization capability of our framework, we also evaluate on an additional dataset which we refer to as ScanNet+ containing additional scenes from ScanNet. Unless otherwise stated, all experiments are performed and averaged over the three standard ScanNet scenes. For evaluation, we use the standard photometric measures PSNR, SSIM, and LPIPS [35] as used in the original NeRF paper [15] and in all baselines we compare to. As we are especially interested in the ability to render accurate depths, we adapt the depth metrics used in DäRF, namely relative error (Rel) and RMSE [25].
We compare our work to several baselines that use depth-supervision for NeRF training in indoor environments. The most related baseline is DäRF [25], as it also leverage a frozen out-of-the-box pre-trained monocular depth prior to supervise NeRF training. SCADE is related to our work as well, but uses a custom-trained depth prior, and does not evaluate the depth accuracy of their method. As SCADE provides pre-trained weights, we evaluate their method on all depth metrics. Additional baselines are a standard NeRF [15], DS-NeRF [7], and DDPrior [22], where we use the results reported by DDPrior. We report results on DDPrior using an out-of-domain prior as trained by SCADE, to replicate our setting where the depth network is trained on a different dataset.
@@ 4.3 Reconstruction Quality
Figure 4: Good RGB rendering quality does not imply good geometric understanding. In this test example, SCADE [26] accurately renders the image (PSNR and SSIM above average), while misinterpreting the geometry of the depicted scene (five-fold average RMSE). The model does not capture the cabinet below the desk in the depth map.
In this section we evaluate our method’s ability to render novel views, and, importantly, its ability to understand the geometric layout of a scene. We thus report both photometric and depth-based metrics. But first, we demonstrate that photometric and depth errors do not always go hand in hand. Observe in Figure 4, that a model can yield great photometric results, while misunderstanding the geometry. In the illustrated example, the cabinet is visible in two training images only and from similar angles. This leads to a misinterpretation of the scene such that the carpet appears to have a beige box pattern. Correcting the geometry in such cases does not improve rendering quality.
@@
RGB-based metrics Depth-based metrics
D PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ SqRel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$
DiffDP [11] - - - 0.100 0.032 0.261 0.128
NeRF [15] 19.03 0.670 0.398 - - 1.163 -
DS-NeRF [7] 20.85 0.713 0.344 - - 0.447 -
DDPrior [22] 20.96 0.737 0.294 - - 0.236 -
DDPrior* [22] 19.29 0.695 0.368 - - 0.474 -
SCADE** [26] 21.54 0.732 0.292 0.086 0.030 0.252 0.118
DäRF [25] 21.58 0.765 0.325 0.151 0.071 0.356 0.168
Ours 21.69 0.737 0.373 0.070 0.024 0.221 0.105
* Trained by SCADE authors. ** Depth metrics evaluated by us based on SCADE’s weights.
@@
We compare our method on RGB-based and depth-based metrics in Table 1. Our method reduces all depth metrics of all baselines by at least $11\%$, and reduces the error of the most related work, DäRF, by $54\%,66\%,38\%$, and $38\%$ on the four depth metrics. Our method also outperforms DDPrior [22] whose depth prior was trained in-domain. Our NeRF setup even improves the initial depth estimates of the depth prior DiffDP. Moreover, our method not only outperforms SCADE, but also avoids the need for a custom-trained depth prior. Training this prior requires evaluating the entire training set 20 times after each epoch to find the best hypothesis for each training example. Our training of SCADE’s cIMLE-based prior takes more than one week on four 24GB NVIDIA Titan RTX GPUs.
When comparing RGB metrics to the closest baseline, SCADE, our method yields a $0.7\%$ worse SSIM, while the PSNR is $0.7\%$ better. Though our method yields an LPIPS that is on average worse than SCADE’s reported result, when reproducing SCADE’s results, some runs lead to an extremely high LPIPS above 0.5 (see Figure 9). Compared to DäRF, our SSIM is $3.7\%$ worse, but all depth metrics are $37\%$ to $66\%$ improved. Overall, we highlight that when geometric consistency is desired, our method enables substantial gains in depth metrics while maintaining comparable RGB metrics.
@@
We qualitatively compare results in Figure 5. In the illustrated examples, our method leads to less severe artifacts, and crisper and more accurate edges in RGB renderings and depth maps. Additional qualitative examples of depth maps are shown in Figure 6. Our model more accurately reconstruct the depicted layouts, like the area underneath the piano in the second image, or the area under the table in the third image. Interestingly, DäRF’s depths are extremely crisp. However, the quantitative evaluation in Table 1 confirms that DäRF learns an inaccurate representation of the geometry. Its strong $L_{2}$-loss enforces the depth maps to look like their prior, although they do not correctly represent the scene.
To demonstrate the robustness of our method we report results on ScanNet+ that were not previously used by the baselines. We observe in Table 2, that our method reduces the RMSE of the baselines by at least $56\%$, while producing comparable photometric results. Please see the supplement for more results.
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$
SCADE [26] 20.57 0.687 0.369 0.396 1.032
DäRF [25] 22.01 0.751 0.319 0.745 1.765
Ours 22.29 0.722 0.391 0.156 0.456